OpenStudy (hayleymeyer):

I love ya 4eve if you help meehh! :( The diameter of Earth is about 7,926 miles and the diameter of Mars is about 4,194 miles. How much more surface area is there on Earth than on Mars to the nearest hundred million square miles? a. 142,000,000 mi2 c. 284,000,000 mi2 b. 14,000,000 mi2 d. 197,000,000 mi

OpenStudy (johnweldon1993):

Hmm, we would call them both Spheres So the surface area would be \(\large SA = 4\pi r^2\) So the surface area of Earth would be \[\large SA_{Earth} = 4 \pi (\frac{7926}{2})^2 = 197,359,487.489mi^2\] \[\large SA_{Mars} = 4 \pi (\frac{4194}{2})^2 = 55,259,471.237mi^2\] So the difference would be \[\large SA_{Earth} - SA_{Mars} = 197,359,487.489mi^2 - 55,259,471.237mi^2\] \[\large SA_{Earth} - SA_{Mars} = 142100016.25mi^2 \] Which is rounded to the nearest hundred million mile of \(\large 142,000,000mi^2\)

OpenStudy (hayleymeyer):

whoa O_O

OpenStudy (hayleymeyer):

thank you!

OpenStudy (johnweldon1993):

Lol yeah...sorry it took so long to type :P But basically find the surface areas of the 2 and subtract em :)

OpenStudy (hayleymeyer):

thanks <33

OpenStudy (johnweldon1993):

Of course :)